The state test scores for 12 randomly selected high school seniors are shown on the right. Complete parts​ (a) through​ (c) below. Assume the population is normally distributed. 1422 1222 985 693 729 840 723 746 542 620 1450 941
Answer #1

Solution:-

(x-u)(x u) 2 (x - u)*2 1422 512.583333 262741.67361 1222 312.583333 97708.340277 985 75.58333335712.8402777 693 -216.41666 46836.173611 729 -180.41666 32550.173611 840 -69.416666 4818.6736111 723-186.41666 34751.173611 746-163.41666 26705.006944 542-367.41666 134995.00694 620 -289.41666 83762.006944 1450 540.583333 292230.34027 941 31.5833333997.50694444 10913 -572.16666 730581.06944 909.41666 -57.216666 73058.106944 Sum Mean

a) The sample mean is 909.417.

b) The sample standard deviation is 257.714.

(n-1)

730581.06944 01 = (12-1)

257.714

c) 95% confidence interval for the population mean is C.I = ( 763.602, 1055.232).

C.I = ar{x}pm z_{alpha /2}ast rac{sigma }{sqrt{n}}

C.I = 909.417 + 1.96*74.3956

C.I = 909.417 + 145.8154

C.I = ( 763.602, 1055.232)